Four different methods of solving a quadratic equation have been discussed in this course: factoring,the square root property,completing the square,and the quadratic formula.Explain under what circumstances each methods would be preferred over any of the other methods. Give an example for each circumstances.
Answer by lwsshak3(11628)

Four different methods of solving a quadratic equation have been discussed in this course: factoring,the square root property,completing the square,and the quadratic formula.Explain under what circumstances each methods would be preferred over any of the other methods. Give an example for each circumstances.
**
quadratic equations are of the standard form: ax^2+bx+c
..
Factoring should be the choice if equation appears to be factorable:
ex: x^2-x+6=0
(x+2)(x-3)=0
x=-2 or x=3
..
Completing the square would be my next choice especially if the coefficient of x^2, a=1, and b & c are small single digits integers:
ex:x^2-4x-6=0
(x^2-4x+4)-6-4=0
(x-2)^2-10=0
(x-2)^2=10
x-2=ħ√10
x=2ħ√10
..
If a, b & c are very large or very small like fractions or decimals use the quadratic formula.
ex: 3x^2+50x-248=0
a=3, b=50, c=-248
x=[-50ħsqrt(50^2-4*3*-248)]2*3
x=[-50ħsqrt(2500+2976)]/6
x=(-50ħ√5476)/6
x=(-50ħ74)/6
x=24/6=4
or
x=-124/6=-62/3
..
I'm not sure what you meant by the square root property.